OpenStudy (anonymous):

WHAT ARE THE SOLUTIONS OF: 2x^2 + 2x - 3?

OpenStudy (amoodarya):

2(x^2+x)-3 =2((x+1/2)^2-(1/2)^2)-3 =2((x+1/2)^2)-2/4-3 =2((x+1/2)^2)-7/2 can you go on ?

OpenStudy (anonymous):

No, I don't know where I would go from there. Are you using the quadratic formula?

OpenStudy (amoodarya):

yes turn it to standard form a(x+h)^2-b=0

OpenStudy (anonymous):

How would I do that?

OpenStudy (amoodarya):

2((x+1/2)^2)-7/2=0 ((x+1/2)^2)=7/4 root of both sides x+1/2=+-root(7/4)